# Pasadena, CA 91125, USA

### v1

@@ 4.3.2 $D_{5}$
from which we can read off the ACSM theory with $U(1)^{6}$ gauge group and 10 chiral multiplets. We find that the superpotential with 8 monopole solutions
$\begin{split}\mathcal{W}&=V_{1}\widetilde{\Phi}_{2}+V_{3}\Phi_{2}\Phi_{6}X+V_{3}\widetilde{\Phi}_{2}\Phi_{6}X+V_{3}\widetilde{\Phi}_{2}\Phi_{4}\Phi_{5}+V_{5}\widetilde{\Phi}_{3}\\ &\quad+V_{2}\Phi_{1}\Phi_{3}+V_{2}\widetilde{\Phi}_{1}\widetilde{\Phi}_{3}+V_{4,6}\widetilde{\Phi}_{3}\end{split}$ (114)
leave the
$U(1)_{A}=U(1)_{T_{2}}+U(1)_{T_{4}}+U(1)_{T_{5}}-U(1)_{\Phi_{1}}-U(1)_{\widetilde{\Phi}_{3}}$ (115)
together with the $U(1)_{f}=U(1)_{T_{6}}-U(1)_{T_{4}}$. We conjecture $U(1)_{A}$ gives the axial $R$ charge under $\mathcal{N}=4$ enhancement. This is tested by computing the $T$-refined superconformal index and its Higgs/Coulomb limits, given as
$\begin{split}\mathcal{I}_{\mathrm{SCI}}(\mathfrak{q},T,z)&=1-\chi_{\mathbf{3}}^{SU(2)}(z)\,T\,\mathfrak{q}^{1/2}+\Big(-1-\chi_{\mathbf{3}}^{SU(2)}(z)+T^{2}\chi_{\mathbf{5}}^{SU(2)}(z)\Big)\mathfrak{q}\\ &\quad-\Big(T^{-1}+T+T^{3}\chi_{\mathbf{7}}^{SU(2)}(z)-T\big(\chi_{\mathbf{3}}^{SU(2)}(z)+\chi_{\mathbf{5}}^{SU(2)}(z)\big)\Big)\mathfrak{q}^{3/2}+\cdots,\\ \mathcal{I}_{C}(\mathfrak{q},z)&=1,\\ \mathcal{I}_{H}(T,z)&=1+\chi_{\mathbf{3}}^{SU(2)}(z)\,T+\chi_{\mathbf{5}}^{SU(2)}(z)\,T^{2}+\chi_{\mathbf{7}}^{SU(2)}(z)\,T^{3}+\cdots.\end{split}$ (116)
The term $\chi^{SU(2)}_{\mathbf{3}}\mathfrak{q}^{\frac{1}{2}}$ in the superconformal index comes from the $SU(2)_{f}$ moment map and also the term $\mathfrak{q}$ corresponds to the $\mathcal{N}=4$ stress-energy tensor. The Coulomb and Higgs limits correctly reproduce the Hilbert series of the trivial Coulomb branch and $\mathbb{C}^{2}/\mathbb{Z}_{2}$ Higgs branch.
The $T$-refined half index matches our proposed expression for the Macdonald index of the $(A_{1},D_{5})$ AD theory, given as
$\begin{split}I\!\!I\!&=\frac{1}{(q)_{\infty}^{6}}\sum_{m\in\mathbb{Z}^{6}}q^{m_{1}m_{2}+m_{2}(m_{2}+m_{4}+m_{5})}q^{\sum_{i=1}^{6}m_{i}}z^{m_{6}-m_{4}}T^{m_{2}+m_{4}+m_{5}}\\ &\quad\times(q^{1+m_{1}}T)_{\infty}(q^{1+m_{1}})_{\infty}(q^{1+m_{2}})_{\infty}^{2}(q^{1+m_{3}}T)_{\infty}(q^{1+m_{3}})_{\infty}(q^{1+m_{4}})_{\infty}(q^{1+m_{5}})_{\infty}\\ &\quad\times(q^{1+m_{6}})_{\infty}(q^{1+m_{4}+m_{5}-m_{6}})_{\infty}\\ &=(q)_{\infty}^{2}(qT)_{\infty}^{2}\sum_{m\in\mathbb{Z}^{6}}\frac{q^{m_{1}m_{2}+m_{2}(m_{2}+m_{4}+m_{5})}q^{\sum_{i=1}^{6}m_{i}}z^{m_{6}-m_{4}}T^{m_{2}+m_{4}+m_{5}}}{(qT)_{m_{1}}(q)_{m_{1}}(q)_{m_{2}}^{2}(qT)_{m_{3}}(q)_{m_{3}}(q)_{m_{4}}(q)_{m_{5}}(q)_{m_{6}}(q)_{m_{4}+m_{5}-m_{6}}}\ .\end{split}$ (117)
Upon series expantion in $q$, we obtain
$\begin{split}I\!\!I\!(q,T,z)&=1+T\chi^{SU(2)}_{\mathbf{3}}(z)q+\left(T+T\chi^{SU(2)}_{\mathbf{3}}+T^{2}\chi^{SU(2)}_{\mathbf{5}}\right)q^{2}+\\ &\quad+\left(T(1+\chi^{SU(2)}_{\mathbf{3}}(z))+T^{2}(\chi^{SU(2)}_{\mathbf{5}}(z)+2\chi^{SU(2)}_{\mathbf{3}}(z))+T^{3}\chi^{SU(2)}_{\mathbf{7}}(z)\right)q^{3}+\cdots,\end{split}$ (118)
@@ 4.4 $(A_{1},D_{2n+2})$ theory
It has an $SU(3)$ flavor symmetr for $n=1$, and $SU(2)\times U(1)$ flavor symmetry for $n>1$. The lattice vectors for the flavor symmetry are given as
@@
We find that this formula is identical to that of the half index of the A-twisted 3d $U(1)^{2n+4}$ Chern-Simons theory with level
@@
The mixed Chern-Simons levels between the gauge $U(1)$s and $R$ symmetry are $\rho=\{2,2,\cdots,2,0,0\}$ for odd $n$ and $\rho=\{1,2,2,\cdots,2,1,0\}$ for even $n$.
@@
Also, we have $4n+4$ chirals multiplets, all with trial $R$ charge 0. Their gauge charges are
@@ 4.4.1 $D_{4}$
There are in total $8$ (dressed) monopole operators of $R=2$ and also neutral under $U(1)_{f}$, listed in Table 5. The global symmetry before turning superpotential is $U(1)^{6}$, and we need to find the superpotential so that we preserve $U(1)_{A}\times U(1)_{f_{1}}\times U(1)_{f_{2}}\subset SO(4)_{R}\times SU(3)$. We found such a superpotential, shown in Table 5. We conjecture that with this superpotential, the theory flows to $\mathcal{N}=4$ SCFT in the IR with enhanced $SU(3)$ flavor symmetry.
@@
where $U(1)_{A}\subset SO(4)_{R}$ and $U(1)_{f_{1}}\times U(1)_{f_{2}}\subset SU(3)$ are the Cartan generators of global symmetry.
Figure 9: The absolute value of the three sphere partition function for the 3d theory of $(A_{1},D_{4})$ as a function of mixing parameter $\nu$ with $U(1)_{A}=U(1)_{\Phi_{2}}$. The function $|Z_{S^{3}}|$ is minimized at $\nu=0$.
To obtain the superconformal $R$ symmetry of this theory, we perform F-maximization. We see that the three-sphere partition function is indeed minimized at $\nu=0$; see Fig. 9.
@@
$\begin{split}\mathcal{I}_{SCI}(\mathfrak{q},T,z)&=1+T\chi^{SU(3)}_{[1,1]}\mathfrak{q}^{\frac{1}{2}}+\left(-1-\chi^{SU(3)}_{[1,1]}+T^{2}\chi^{SU(3)}_{[2,2]}\right)\mathfrak{q}\\ &\quad+\left(T(1-\chi^{SU(3)}_{[2,2]}-\chi^{SU(3)}_{[3,0]}-\chi^{SU(3)}_{[0,3]})+T^{3}\chi^{SU(3)}_{[3,3]})\right)\mathfrak{q}^{\frac{3}{2}}+\cdots,\\ \mathcal{I}_{C}(\mathfrak{q},z)&=1,\\ \mathcal{I}_{H}(T,z)&=1+\chi_{[1,1]}^{SU(3)}T+\chi_{[2,2]}^{SU(3)}T^{2}+\chi_{[3,3]}^{SU(3)}T^{3}+\chi_{[4,4]}^{SU(3)}T^{4}+\cdots\ ,\end{split}$ (135)
@@
The $T$-refined half index matches our conjectured formula for the Macdonald index of $(A_{1},D_{4})$ AD theory:
@@ 4.4.2 $D_{6}$
Before turning on the superpotential, this theory has $U(1)^{8}$ flavor symmetry. We look for an appropriate superpotential that breaks all the symmetry except for $U(1)_{A}\times U(1)_{f_{1}}\times U(1)_{f_{2}}$. Since this theory has a rather high rank, we were not able to search through all possible monopole combinations for superpotential deformation. But we do find a good candidate superpotential given as
@@
We conjecture that upon turning on this superpotential, the ACSM theory flows to $\mathcal{N}=4$ SCFT in the IR with $SU(2)\times U(1)$ flavor symmetry. We identify the axial symmetry $U(1)_{A}\subset SO(4)_{R}$ is identified with
@@
Since this theory has a rather high rank, we were not able to find a superconformal index to a reasonable order to test the SUSY enhancement, and the Coulomb/Higgs moduli space. However, we do find that the $U(1)_{A}$ refined half index of our theory indeed reproduces the Macdonald index of $(A_{1},D_{6})$ theory:
$\begin{split}I\!\!I\!(&q,T,w,x)\\ &=\frac{1}{(q)_{\infty}^{8}}\sum_{m\in\mathbb{Z}^{8}}q^{\frac{1}{2}m^{t}b^{D_{6}}m}(q^{\frac{1}{2}})^{m_{1}+m_{7}+2\sum_{i=2}^{6}m_{i}}(w^{2})^{m_{1}-m_{5}-m_{7}+m_{8}}\left(-\frac{x}{w}\right)^{m_{1}-m_{7}}T^{m_{3}+m_{5}+m_{6}+\frac{1}{2}(m_{1}+m_{7})}\\ &\quad\times\left(\prod_{i=1}^{8}(q^{1+m_{i}})_{\infty}\right)(Tq^{1+m_{2}})_{\infty}(Tq^{1+m_{4}})_{\infty}(q^{1+m_{1}+m_{3}-m_{7}})_{\infty}(q^{1-m_{1}+m_{5}+m_{6}+m_{7}-m_{8}})_{\infty}\\ &=(q)_{\infty}^{2}(qT)_{\infty}^{2}\\ &\times\sum_{m\in\mathbb{Z}^{8}}\frac{q^{\frac{1}{2}m^{t}b^{D_{6}}m}(q^{\frac{1}{2}})^{m_{1}+m_{7}+2\sum_{i=2}^{6}m_{i}}(w^{2})^{m_{1}-m_{5}-m_{7}+m_{8}}(\frac{x}{w})^{m_{1}-m_{7}}T^{m_{3}+m_{5}+m_{6}+\frac{1}{2}(m_{1}+m_{7})}}{\left(\prod_{i=1}^{8}(q)_{m_{i}}\right)\left(\prod_{j=2,4}(qT)_{m_{j}}\right)(q)_{m_{1}+m_{3}-m_{7}}(q)_{-m_{1}+m_{5}+m_{6}+m_{7}-m_{8}}}\end{split}$ (142)
Indeed the half index reproduces our conjectured fermionic sum formula. Upon series expansion, we obtain
$\begin{split}I\!\!I\!(q,T,w,x)&=1+qT\left(w^{2}+\frac{1}{w^{2}}+2\right)+q^{3/2}T^{3/2}\left(wx+\frac{w}{x}+\frac{1}{wx}+\frac{x}{w}\right)\\ &+q^{2}\left(T^{2}\left(w^{4}+\frac{1}{w^{4}}+2w^{2}+\frac{2}{w^{2}}+3\right)+T\left(w^{2}+\frac{1}{w^{2}}+3\right)\right)\\ &+q^{5/2}\Bigg(T^{5/2}\left(w^{3}x+\frac{w^{3}}{x}+\frac{x}{w^{3}}+\frac{1}{w^{3}x}+2wx+\frac{2w}{x}+\frac{2}{wx}+\frac{2x}{w}\right)\\ &\qquad+T^{3/2}\left(wx+\frac{w}{x}+\frac{1}{wx}+\frac{x}{w}\right)\Bigg)+\cdots\ ,\end{split}$ (143)

### v2

@@ 4.3.2 $D_{5}$
from which we can read off the ACSM theory with $U(1)^{6}$ gauge group and 10 chiral multiplets. We conjecture that there exists a superpotential deformation with 8 monopole operators
$\begin{split}\mathcal{W}&=V_{1}\widetilde{\Phi}_{2}+V_{3}\Phi_{2}\Phi_{6}X+V_{3}\widetilde{\Phi}_{2}\Phi_{6}X+V_{3}\widetilde{\Phi}_{2}\Phi_{4}\Phi_{5}+V_{5}\widetilde{\Phi}_{3}\\ &\qquad+V_{2}\Phi_{1}\Phi_{3}+V_{2}\widetilde{\Phi}_{1}\widetilde{\Phi}_{3}+V_{4,6}\widetilde{\Phi}_{3}\end{split}$ (114)
leaving the following symmetry unbroken
$U(1)_{A}=U(1)_{T_{2}}+U(1)_{T_{4}}+U(1)_{T_{5}}-U(1)_{\Phi_{1}}-U(1)_{\widetilde{\Phi}_{3}}\ ,$ (115)
together with $U(1)_{f}=U(1)_{T_{6}}-U(1)_{T_{4}}$. As before, this proporsal is tested against the $T$-refined superconformal index and its Higgs/Coulomb limits, given as
$\begin{split}\mathcal{I}_{\mathrm{SCI}}(\mathfrak{q},T,z)&=1-\chi_{\mathbf{3}}^{SU(2)}(z)\,T\,\mathfrak{q}^{1/2}+\Big(-1-\chi_{\mathbf{3}}^{SU(2)}(z)+T^{2}\chi_{\mathbf{5}}^{SU(2)}(z)\Big)\mathfrak{q}\\ &\qquad-\Big(T^{-1}+T+T^{3}\chi_{\mathbf{7}}^{SU(2)}(z)-T\big(\chi_{\mathbf{3}}^{SU(2)}(z)+\chi_{\mathbf{5}}^{SU(2)}(z)\big)\Big)\mathfrak{q}^{3/2}+\cdots,\\ \mathcal{I}_{C}(\mathfrak{q},z)&=1,\\ \mathcal{I}_{H}(T,z)&=1+\chi_{\mathbf{3}}^{SU(2)}(z)\,T+\chi_{\mathbf{5}}^{SU(2)}(z)\,T^{2}+\chi_{\mathbf{7}}^{SU(2)}(z)\,T^{3}+\cdots.\end{split}$ (116)
The term $\chi^{SU(2)}_{\mathbf{3}}\mathfrak{q}^{\frac{1}{2}}$ in the superconformal index comes from the $SU(2)_{f}$ moment map and the term $\mathfrak{q}$ corresponds to the $\mathcal{N}=4$ stress-energy tensor. The Coulomb and Higgs limits correctly reproduce the Hilbert series of the trivial Coulomb branch and $\mathbb{C}^{2}/\mathbb{Z}_{2}$ Higgs branch.
The $T$-refined half-index matches our proposed expression for the Macdonald index of the $(A_{1},D_{5})$ AD theory, given as
$\begin{split}I\!\!I\!&=\frac{1}{(q)_{\infty}^{6}}\sum_{m\in\mathbb{Z}^{6}}q^{m_{1}m_{2}+m_{2}(m_{2}+m_{4}+m_{5})}q^{\sum_{i=1}^{6}m_{i}}z^{m_{6}-m_{4}}T^{m_{2}+m_{4}+m_{5}}\\ &\qquad\times(q^{1+m_{1}}T)_{\infty}(q^{1+m_{1}})_{\infty}(q^{1+m_{2}})_{\infty}^{2}(q^{1+m_{3}}T)_{\infty}(q^{1+m_{3}})_{\infty}(q^{1+m_{4}})_{\infty}\\ &\qquad\times(q^{1+m_{5}})_{\infty}(q^{1+m_{6}})_{\infty}(q^{1+m_{4}+m_{5}-m_{6}})_{\infty}\\ &=(q)_{\infty}^{2}(qT)_{\infty}^{2}\sum_{m\in\mathbb{Z}^{6}}\frac{q^{m_{1}m_{2}+m_{2}(m_{2}+m_{4}+m_{5})}q^{\sum_{i=1}^{6}m_{i}}z^{m_{6}-m_{4}}T^{m_{2}+m_{4}+m_{5}}}{(qT)_{m_{1}}(q)_{m_{1}}(q)_{m_{2}}^{2}(qT)_{m_{3}}(q)_{m_{3}}(q)_{m_{4}}(q)_{m_{5}}(q)_{m_{6}}(q)_{m_{4}+m_{5}-m_{6}}}\ .\end{split}$ (117)
Upon series expansion in $q$, we obtain
$\begin{split}I\!\!I\!(q,T,z)&=1+T\chi^{SU(2)}_{\mathbf{3}}(z)q+\left(T+T\chi^{SU(2)}_{\mathbf{3}}+T^{2}\chi^{SU(2)}_{\mathbf{5}}\right)q^{2}+\\ &\qquad+\left(T(1+\chi^{SU(2)}_{\mathbf{3}}(z))+T^{2}(\chi^{SU(2)}_{\mathbf{5}}(z)+2\chi^{SU(2)}_{\mathbf{3}}(z))+T^{3}\chi^{SU(2)}_{\mathbf{7}}(z)\right)q^{3}+\cdots,\end{split}$ (118)
@@ 4.4 $(A_{1},D_{2n+2})$ theory
It has an $SU(3)$ flavor symmetry for $n=1$, and $SU(2)\times U(1)$ flavor symmetry for $n>1$. The lattice vectors for the flavor symmetry are given as
@@
Again, we start from an observation that this formula is identical to the half-index of the 3d $U(1)^{2n+4}$ Chern-Simons matter theory with level
@@
The mixed Chern-Simons levels between the gauge $U(1)$s and $R$-symmetry are $\rho=\{2,2,\cdots,2,0,0\}$ for odd $n$ and $\rho=\{1,2,2,\cdots,2,1,0\}$ for even $n$.
@@
Also, we have $4n+4$ chiral multiplets, all with trial $R$-charge 0. Their gauge charges are
@@ 4.4.1 $D_{4}$
There are $8$ dressed monopole operators in total, with $R=2$ and neutral under $U(1)_{f}$, listed in Table 5. The global symmetry before turning on the superpotential is $U(1)^{6}$, and we expect that there exists a superpotential deformation that preserves $U(1)_{A}\times U(1)_{f_{1}}\times U(1)_{f_{2}}\subset SO(4)_{R}\times SU(3)$. We find that the superpotential shown in Table 5 satisfies this requirement. We conjecture that with this superpotential, the theory flows to $\mathcal{N}=4$ SCFT in the IR with enhanced $SU(3)$ flavor symmetry.
@@
where $U(1)_{A}\subset SO(4)_{R}$ and $U(1)_{f_{1}}\times U(1)_{f_{2}}\subset SU(3)$ are the Cartans of global symmetry.
Figure 9: The absolute value of the three-sphere partition function for the 3d theory of $(A_{1},D_{4})$ as a function of the mixing parameter $\nu$ with $U(1)_{A}=U(1)_{\Phi_{2}}$. The function $|Z_{S^{3}}|$ is minimized at $\nu=0$.
To obtain the superconformal $R$-symmetry of this theory, we perform F-maximization. We see that the three-sphere partition function is indeed minimized at $\nu=0$; see Figure 9.
@@
$\begin{split}\mathcal{I}_{SCI}(\mathfrak{q},T,z)&=1+T\chi^{SU(3)}_{[1,1]}\mathfrak{q}^{\frac{1}{2}}+\left(-1-\chi^{SU(3)}_{[1,1]}+T^{2}\chi^{SU(3)}_{[2,2]}\right)\mathfrak{q}\\ &\qquad+\left(T(1-\chi^{SU(3)}_{[2,2]}-\chi^{SU(3)}_{[3,0]}-\chi^{SU(3)}_{[0,3]})+T^{3}\chi^{SU(3)}_{[3,3]})\right)\mathfrak{q}^{\frac{3}{2}}+\cdots,\\ \mathcal{I}_{C}(\mathfrak{q},z)&=1,\\ \mathcal{I}_{H}(T,z)&=1+\chi_{[1,1]}^{SU(3)}T+\chi_{[2,2]}^{SU(3)}T^{2}+\chi_{[3,3]}^{SU(3)}T^{3}+\chi_{[4,4]}^{SU(3)}T^{4}+\cdots\ ,\end{split}$ (135)
@@
The $T$-refined half-index in the A-twisted limit matches our conjectured formula for the Macdonald index of $(A_{1},D_{4})$ AD theory:
@@ 4.4.2 $D_{6}$
from which we can read off the ACSM theory with $U(1)^{8}$ gauge group and 12 chiral multiplets.
Before turning on the superpotential, this theory has $U(1)^{12}$ flavor symmetry. We look for an appropriate superpotential that breaks all the symmetry except for $U(1)_{A}\times U(1)_{f_{1}}\times U(1)_{f_{2}}$. Since this theory has a rather high rank, we were not able to search over all possible monopole combinations for superpotential deformation. But we do find a good candidate superpotential given as
@@
We conjecture that upon turning on this superpotential, the ACSM theory flows to $\mathcal{N}=4$ SCFT in the IR with $SU(2)\times U(1)$ flavor symmetry. We identify the axial symmetry $U(1)_{A}\subset SO(4)_{R}$ as
@@
Since this theory has a rather high rank, we were not able to find a superconformal index to a reasonable order to test the SUSY enhancement, and the Coulomb/Higgs moduli space. However, we do find that the $U(1)_{A}$ refined half-index of our theory indeed reproduces the Macdonald index of $(A_{1},D_{6})$ theory:
$\begin{split}I\!\!I\!(&q,T,w,x)\\ &=\frac{1}{(q)_{\infty}^{8}}\sum_{m\in\mathbb{Z}^{8}}q^{\frac{1}{2}m^{t}b^{D_{6}}m}(q^{\frac{1}{2}})^{m_{1}+m_{7}+2\sum_{i=2}^{6}m_{i}}(w^{2})^{m_{1}-m_{5}-m_{7}+m_{8}}\left(\frac{x}{w}\right)^{m_{1}-m_{7}}T^{m_{3}+m_{5}+m_{6}+\frac{1}{2}(m_{1}+m_{7})}\\ &\quad\times\left(\prod_{i=1}^{8}(q^{1+m_{i}})_{\infty}\right)(Tq^{1+m_{2}})_{\infty}(Tq^{1+m_{4}})_{\infty}(q^{1+m_{1}+m_{3}-m_{7}})_{\infty}(q^{1-m_{1}+m_{5}+m_{6}+m_{7}-m_{8}})_{\infty}\\ &=(q)_{\infty}^{2}(qT)_{\infty}^{2}\\ &\times\sum_{m\in\mathbb{Z}^{8}}\frac{q^{\frac{1}{2}m^{t}b^{D_{6}}m}(q^{\frac{1}{2}})^{m_{1}+m_{7}+2\sum_{i=2}^{6}m_{i}}(w^{2})^{m_{1}-m_{5}-m_{7}+m_{8}}(\frac{x}{w})^{m_{1}-m_{7}}T^{m_{3}+m_{5}+m_{6}+\frac{1}{2}(m_{1}+m_{7})}}{\left(\prod_{i=1}^{8}(q)_{m_{i}}\right)\left(\prod_{j=2,4}(qT)_{m_{j}}\right)(q)_{m_{1}+m_{3}-m_{7}}(q)_{-m_{1}+m_{5}+m_{6}+m_{7}-m_{8}}}\end{split}$ (142)
Indeed, the half-index reproduces our conjectured fermionic sum formula. Upon series expansion, we obtain
$\begin{split}I\!\!I\!(q,T,w,x)&=1+qT\left(w^{2}+\frac{1}{w^{2}}+2\right)+q^{3/2}T^{3/2}\left(wx+\frac{w}{x}+\frac{1}{wx}+\frac{x}{w}\right)\\ &\qquad+q^{2}\left(T^{2}\left(w^{4}+\frac{1}{w^{4}}+2w^{2}+\frac{2}{w^{2}}+3\right)+T\left(w^{2}+\frac{1}{w^{2}}+3\right)\right)\\ &\qquad+q^{5/2}\Bigg(T^{5/2}\left(w^{3}x+\frac{w^{3}}{x}+\frac{x}{w^{3}}+\frac{1}{w^{3}x}+2wx+\frac{2w}{x}+\frac{2}{wx}+\frac{2x}{w}\right)\\ &\qquad\qquad\qquad+T^{3/2}\left(wx+\frac{w}{x}+\frac{1}{wx}+\frac{x}{w}\right)\Bigg)+\cdots\ ,\end{split}$ (143)